Protected:

This content is password protected. To view it please enter your password below:

Posted in Lab Meeting | Comments Off on Protected:

Thursday 08/27/15

9:30 am – 7:00 pm,

Lab meeting prep

  • working on slides for Group B presentation, 9:30 am – 1:30 pm

Two-color data analysis

  • interesting results from ANT-C G6 quantification
  • median overlap ~25%, median entanglement ~0.06
  • need to make figures still

Meetings

  • requested to make MERFISH slides for grant meeting
  • practice talk to be given Sept 4 (next Friday)

Sequencing prep

  • ran tape-station on 1:100 dilution samples
  • looks good, except I cant’ connect to the data folder: https://data.rc.fas.harvard.edu
  • 600 nM * (450/300) = 900 nM starting. Desired 10 nM. So my 1:100 dilutions are around 9 nM, pretty good.
  • submitted sequencing request (still did not submit physical samples, should follow up with Christian).

To do

  • reads per domain for all domains (by color) compared to green domains
  • analyze the BX-C F6 data
  • images of the BX-C F6 data for presentation
  • quantification comparison of BX-C F6

For later

  • restain Ph with shorter fix time
  • try Ph stain on MeOH fixed cells
Posted in Summaries | Comments Off on Thursday 08/27/15

Wednesday 08/26/19

9:30 am – 9:00 pm

Goals

  • ChromatinCropper2 two color data for ANT-C G6
  • Spot finding for two color BX-C F6 data
  • ChromatinCropper2 two color data
  • tape station of sequencing prepped library for Ph-KD and control
  • qPCR of sequencing prepped library for Ph-KD and control

STORM

  • stop STORM run, clean up STORM2 for Yari
  • analyzing data on STORM2
    • shouldn’t launch more than 30 jobs at once due to read/write limitations through the USB3.0 port on Morgan — makes interacting with the disk impossible.

Data Analysis

  • troubleshooting issues with chromatic alignment
  • fixed issues with chromatic alignment
    FixedAlignmentBug WarpErrorAllDownAndLeft
  • Cropper2 not applying warps properly: previous warpfile was still hardcoded, not being passed
  • fixed warp for alignment of conventional images.
  • still need to make sure warp gets applied to STORM images.

Deep seq prep

  • rerunning Kappa quantification
  • kappa standards 1:6 in triplicate (A1-6 through B1-6, C1-6) + No template controls A7-C7
  • original dilution 1:100
  • serial dilution (in triplicate)
  • 1:100,000, 1:1,000,000, 1:10,000,000
  • 10 in 990.
  • just used median in fit. linFit(x) = p1*x + p2; p1 = -0.3005 (-0.3083, -0.2927), p2 = 4.989 (4.823, 5.156)
  • looks good
    KapaQuantPlate_150826 KapaQuant_150826
  • 1 in 10,000 dilution is ~60 pM. so original is 600 nM, assuming average molecule size of 452. Otherwise correct by multipling by 452/aveSize.

Genome analysis

  • figured out header structure and wrote missing header for modENCODE Kc167 SAM data.
  • saved as dm3_SAM_header.txt on Morgan on \\MorganData1\Chromatin\2015-08-25_KcENCODEdata
Posted in Summaries | Comments Off on Wednesday 08/26/19

Tuesday 08/25/15

9:30 am – 10:15 pm

STORM

  • finish overnight run
  • switching still looks decent
  • setup data to analyze on MORGAN.
  • updated z-calibrations for both 647 and 750 parameters
  • fitting bead data for drift correction
  • all jobs launched before 1pm (probably need another 2 hours to finish running)

New experiments

  • stained BX-C and F6 (using previous mixed sample)
  • denatured at upstairs heat block at 78 (temp reasonably sensitive to adding water, use with care)
  • hybridized from 10am – 7pm at 47C
  • rinsed and imaging in MEA and freshly made Glox (left glox made yesterday out last night, tossed it)
  • spots switching much better in 750 (Ben also fixed the TIRF)
  • collecting ~55K frames 750 (30Hz, camera won’t do faster) and 65K frames 650 overnight

Ph-project

  • assemble figures, send to NF
  • latest complete version sent to all authors

Analysis of Green data

  • mapping total reads per region
  • Morgan connection got really slow

counting total reads per region

I used to use a “GenerateStrandCount” function you had written (I believe in matlab-functions).
I presume this has been replaced by “GenerateCoverageVector” or “GenerateSequenceDataStructure”?
Both of these try to find a “SequenceDictionary” field. The structure I get from BioIndexedFile on my SAM file doesn’t have this field.
I believe this is because my SAM file has no header. That pissed off cufflinks too, though I eventually got it to run. modENCODE doesn’t seem to believe in headers.

brute force approach

  • using fseek to count reads, no fancy coverage vector for whole genome
  • running over night

To do

  • format figures into combined PDF for Ph-Polymerization paper
  • Analyze Hi-C data
  • Organize presentation of Green data
  • Analyze two-color data
Posted in Summaries | Comments Off on Tuesday 08/25/15

Monday 08/24/15

9:30 am – 5:15 pm, 8:20 pm – 11:30 pm

Ph-polymerization

Manuscript to do

  1. STORM section, reply to notes in main text
  2. Finish Layout Supp figure 2 and Supp figure 4
    • added supplement captions form both
  3. Check the legend to Figure 5 (new model) and sup. Fig. 11 (rest of model fig)
  4. Compile everything into a PDF
  5. Check response to reviewers.

STORM

  • imaging on STORM2
  • 2 color
  • MEA COT buffer
  • sample quality (ANTC-750 + G6-647)
    • still an odd background nuclear glow in 647 channel (not related to STORM-dye switching, uniform, unbleaching. Still suspect some contaminate Draq5 or something of the sort. Maybe the formamide overnight treatmetn helped drop this down, the STORM switching on top of this background looks pretty good.
  • 561 beads just switch dark / bleach quickly in the MEA COT buffer
    • swapped fresh buffer, added fresh beads. Conventional imaging beads nice and bright
    • start STORM imaging, and beads go dark pretty quick and don’t come back
    • I recall this happening before, and only lasting the first few movies
    • indeed the beads do come back bright, but now the 750 switching looks bad.
    • probably the buffer quality is decreasing, favoring the beads and disfavoring the dyes again. Unfortunately the switching quality in 750 is a no-go at this point.
    • I want to try this again in BME
  • switched buffer to BME.
    • beads don’t recover very well. They aren’t as completely dim but its still pretty bad. Should replace with fresh beads again.
    • 750 switching not as bright, and most dramatically, not as switchy — fewer blinking dyes for way less long. Need to recalibrate all my ramp files
    • using substantial 405 I can get some switching. Let’s see how this goes
    • next I’ll try MEA alone, no COT.
  • switched to MEA no COT
    • first movie, 750 dim but blinking, 45K frames instead of 70 (better than the BME, but we’ll see if it lasts)
    • at least the newly added beads stay bright
    • running the MEA overnight.
    • still a total unusual amount of background in the 647 channel

Cell culture

  • Psc null cells growing well. I have 3 vials in freezer so hopefully those are fine if I need them again.
  • Terminated Psc null flasks
  • trimmed down to 1 KC in SFX flask and 1 KC in Schneider’s flask.
  • interestingly the Schneider’s cells are now adhering better than the SFX

Seq Prep

  • Tape station training
  • tape needs to be reserved online
  • ordered 1 tape for next set of experiments
  • HS DNA is the recommended tape for library prep.
  • protocols sent in email.
Posted in Summaries | Comments Off on Monday 08/24/15

Friday 08/21/15

9:30 am – 7:00 pm

Ph project

  • rereading and revising manuscript
  • sent comments back to team
  • revising layout of Fig 5.
  • revising response to reviewers
  • sent updates back to team

STORM imaging

  • Jeff realigned lasers on STORM4 table, fixing the 90% power loss on the 647 laser due to poor fiber coupling.
  • quadview insert and master need to match. Swapping these may have damaged the mirror on the quadview for the 560 channel on STORM4 (run with STORM2’s insert) and mirror for the quadview on STORM3 (running with STORM3’s insert).
  • can probably run with shutters…

Chromatin analysis

  • ChromatinCropper is crashing at the scatter command for large fields of view
  • fixed, change to plot
  • analyzing Green domains
  • forgot to change npp back to 160, will do in post analysis
Posted in Summaries | Comments Off on Friday 08/21/15

Thursday 08/20/15

9:30 am – 1:30 am

Microscope troubleshooting

  • might be able to try 2-color imaging in STORM5 with George E and shutters.
  • doesn’t work. No emission filter to block 405
  • 2 color conventional images of my stains look great though.
  • trying to image on STORM4
  • issues with quadview – 561 channel not working

<

del datetime=”2015-08-21T21:53:51+00:00″>

Writing

  • revise draft of Ph-manuscript and send replies to Nicole and Ajaz

RNAi experiments

  • Day 4 for KD. Lyse cells, perform RNA extraction.
    • collected 7 tubes of cells, concentrated in spins
    • just froze cells in lyses buffer at -80C. Will continue these when I have time.
  • fix some cells for imaging backup
    • prep cover slips
    • plate cells, allow 3 hours to attach
    • fixed 4 coverslips of Ph KD and 4 coverslips of mock.
    • these may be necessary for continuing the multicolor experiments

Meetings

  • discussion with XZ about grant, 11am
  • update on revision experiments, 1pm

Live imaging

  • TIRF2-mmaple on STORM4
    • sample don’t look promising
    • snapped some images, saved on STORM4, need to copy over for update meeting
  • TIRF2-GFP on STORM3 (Bogdan, 6pm)
    • samples not well adhered (polylysine treatment was omitted).
Posted in Summaries | Comments Off on Thursday 08/20/15

Wednesday 08/19/15

9:00 am – 11:30 pm

Visiting postdoc candidate interviews

  • presentation 11 – 1:30 pm
  • lunch 1:30 – 2:00 pm
  • individual meeting 4-4:30 pm

STORM

  • finish imaging 2 color
  • take bead images for z-calibration
  • take bead images for chromatic correction

Chromatin analysis

  • selecting images of internal domains for presentation for XZ
  • Analyzing SIM data. Images look on par with confocal images. Registration looks similar. Cell numbers are lower.
  • processing the 2-color data for BXC and L2 in Ph-KD
  • updating ChromatinCropper2 GUI to handle multicolor
  • Analyzing 2-color data
    • almost all the 750 spots disappeared in both conventional and STORM images.
    • I swear these were visible earlier.
    • I think maybe its a focal plane issue, somehow the z-offsets are larger now.
    • Maybe Hal’ interpretation of the direction of the offset has also changed, +/- relative to wavelength is interpreted differently on the different systems.

Presentations

  • starting slides for meeting with XZ tomorrow
  • touch up slides for discussion with post-doc candidate
    • not so many cartoons of chromatin, here words will actually be better.
Posted in Summaries | Comments Off on Wednesday 08/19/15

Tuesday 08/18/15

9:45 am – 8:00 pm,
remotely, 9:15 pm – 11:00 pm

RNAi validation

Finish library prep

  • run final bead purification and elution

background prep

  • contact Bauer core to confirm dates for tape analyzer
  • reading background for KAPA_Library_Quantification
    • stresses don’t freeze thaw repeatedly dilute DNA samples
    • keep dilute DNA samples in a buffer, otherwise they degrade rapidly
    • avoid multichannel pipettes and multiwell plates, but run lots of replicates (not convienent)
  • reading protocol for KAPA Library Quantification kit
    • got Kapa master mix from Jeff (has primers added)

setup

  • 6 standards + NT control, 3 replicates (21)
  • library dilutions (3 libraries)
    • dilute each library 1:100. Lets do 2 library + 198 buffer. These are the working dilutions (200 uL).
    • target dilutions: 1:800; 1:3,200; 1:128,000
    • 10 uL working-dilution + 70 uL buffer (1:8 of a 1:100). total 80 uL
    • 20 uL previous-dilution + 60 uL buffer (1:4 of 1:800). total 80 uL
    • 20 uL previous-dilution + 60 uL buffer (1:4 of 1:3200). total 80 uL
  • do this twice from the same working-dilution for each library (2 reps x 3 dils x 3 libs) = 18
  • setup in 96 well plate (39 samples, 3×7 and 3x(3*2)).
  • all dilutions still above the concentration range of the standards. Should repeat setup.
    QPCR_KapaQuantResults_1
  • some pipetting inaccuracies also detected.

Imaging

  • staining Telomere-GFP samples with Draq5
  • added Draq5 to Ph-KD and mock-KD cells stained previously with Ph-Cy3B
  • quantify samples on confocal LSM700 (11am – 1pm)
  • stained Ph-KD and mock-KD with Hoechst. Looks better than Draq5 on fixed cells.

Editing

  • finish comments for Ye. Send back.

Coding: Quantify Fixation GUI

  • adding alignment stats to GUI step 2

Data analysis

  • analyzing internal data for BX-C L4E22to24
Posted in Summaries | Comments Off on Tuesday 08/18/15

Monday 08/17/15

9:45 am

RNA library prep

setup

  • protocol from NEB
  • discuss plans with Jeff
  • request training on BioAnalyzer and/or TapeStation instrument for library QC
  • collecting kit components

RNA purification

  • run out 1 uL on a 1% agarose gel in TAE
  • treat samples with NEB DNase I (1 uL in 20 uL reaction, used RQ1 DNase buffer, don’t have an NEB one. validated the RQ1 is also a DNase I from Promega, guaranteed RNase free)
  • Purify samples on AMPure RNA XP beads.
  • performed all washes on column
  • gel results: (clearly degraded). RNA is a month old (stored at -80C, but used for several RTPCR runs.
    degraded_TotRNA
  • probably do better to start 1 clean fresh run with newly isolated samples.
  • we can run through the procedure today with these samples to get familiar with the protocol.

First strand synthesis

  • ran without actinomyocin (don’t have any). Already did DNA digestion so hopefully this isn’t necessary

Second strand synthesis

  • setup is easy
  • run bead cleanup using Aline beads from Jeff in place of AMPure XP beads

More steps

  • purify dsDNA (beads)
  • end prep
  • adapter ligation
  • purify DNA (beads)
    • this uses a lot of beads. Ordered more Aline beads from Aline biosciences
  • PCR enrichment
    • ran on QPCR machine with EvaGreen
    • 14 cycles, started to saturate, pulled off
    • testing sample by running out product on a gel
    • Typhoon isn’t working — images are perfectly 100% blank. no noise, no contrast, no background of gel.
  • still have 1 more bead clean up to do.

STORM analysis

  • compute z-calibration from surface beads for STORM2 for Ella
  • cancelled STORM imaging for tonight (insufficient time to setup a run before midnight, insufficient time to collect 2 color data if the staining works before 10am).
  • signed up for STORM4 for Thursday and Friday for live imaging with Steven. STORM4 might not work — the 100x objective went missing.

Ph KD immuno

  • finished staining
  • tested samples on Turnkey. WT shows decent staining. Ph in KD not detectable, but need nuclear counterstain to validate cell region.
  • maybe try imaging these on confocal tomorrow.

Editing

  • Editing recommendations for YF application
  • Editing Ph manuscript draft
  • got new Ph manuscript draft, will try to add to current one.
Posted in Summaries | Comments Off on Monday 08/17/15